OpenStudy (anonymous):

how do i solve: x+4t=4

Parth (parthkohli):

Solve for?

OpenStudy (anonymous):

x

Parth (parthkohli):

Subtract 4t from both sides... that's it!

Parth (parthkohli):

What you do in solving for a variable is that you get the variable by itself, okay?

OpenStudy (anonymous):

here we have two variables, so we NEED at least 2 equations to solve this system... but if your just wanting to solve for x then it would be x=4-4t

OpenStudy (anonymous):

and so how do i solve this one? 2x-2=t+6

Parth (parthkohli):

For \(x\)? You have to add 2 to both sides, then divide both by 2.

OpenStudy (anonymous):

so this is what it looks like after adding 2 to both sides. 2x= t+8

OpenStudy (anonymous):

put t=0 then solve.

OpenStudy (anonymous):

like, what do you mean t=0?

OpenStudy (anonymous):

x+4*0=4 x=4

OpenStudy (anonymous):

ok, kinda confused now.

OpenStudy (anonymous):

the problem was 2x-2=t+6, i added 2 to both sides and got 2x=t+6 what happens next?

OpenStudy (anonymous):

value of x?

OpenStudy (anonymous):

yea, im looking for the value of x.

OpenStudy (unklerhaukus):

2x-2=t+6 2x-2+2=t+6+2 2x=t+6+2 x=(t+6+2)/2

OpenStudy (unklerhaukus):

you can simplify further

OpenStudy (anonymous):

ok, thank you so much!!!! :)

OpenStudy (unklerhaukus):

what your simplified solution for x?